20140615 D(+some E) night with LBCs/MODS1/LUCI1/LBCs

Some cirrus before sunset. LBCs OK, MODS1 OK. LUCI1 noisy, but not so bad as last night.

Plan for 20140615

LBCs for Saturn --- start focus/collimate about 8:45pm
   * for a total of about 20 min open shutter time,
      * if sky @ r < 12,000 cnts, 2 x OB10 (9 x 60-sec) or
      * if sky @ r > 12,000 cnts, 4 x OB20 (9 x 30-sec)
Switch to MODS MODS at about 10pm:
   * target 1 (~1 hr)
      * J1324_1asec_HA2to3.5.acq : R=13.89 star, but near -Y lim of patrol field, or R=15.6 star closer to axis
      * J1324_1asec.obs : 600sec x 3 (will repeat for one more 600-sec exposure). R=17.18
      * HIP66935
   * target 2 (~1 hr)
      * J1435_1asec_HA1to3.acq : 
      * J1435_1asec.obs : 600 x 3 g=16.8
      * HIP71172 
   * spectrophot standard in dual mode 
      * BD+33 2642 
   * target 3: 
      * PDS456 K1 for 1 hour (close to moon...) --- 
         * PDS456k1.acq
         * PDS456_600sec.obs (which will offset by 2" along slit) i
       * if moon too bright, could try another bright quasar... or switch directly to LUCI1
Switch to LUCI1 (1:00 am):
   * LBV star
   * Doug's TT tests
Switch to LBCs at 2am
   * OB11 5 x (or run OB11_5 with will do a 5-pt dither)
   * OB12 to check background counts
      * if OK, loop 10 times over OB12,OB13,OB14 (use /home/lbcobs/supportscripts/SymbioticSupport/loop_observe234.py)
   * finish with repeats on OB11 (use loop_observe11.py in same directory)

Summary

A night of inefficiencies and some weather problems, though some data were taken and new things learned.
  • Started with LBCs to observe field. Tried dowrs.
  • Moved to MODS and observed 1 quasar and cals.
  • Switched to LUCI1 and observed the LBV star and cal after some initial observer errors. LUCI1 is looking OK, though there is still pickup noise visible in many channels, not just limited to the two in the upper right quadrant. But, unlike last night, the darks and low-background fields showed some positive counts in all channels, not zeros.
  • Back to the LBCs to finish the night.

Log

02:55 Open & slew to MABlank9 for g' and r' sky flats. Still too bright.

03:05 Starting twilight flats. 15-10k at g' and 9-6k at r' with exptime 1:2.2 sec for g':r'

03:12 RPC error on Blue camera. CCDTemp was wacky, something like 10 digits! As the error happened right after the first OB finished (lost last Blue image), tried to load & play the next (180 = PA) OB, but got the message about trying to recover after a crash. LBC recovered OK.

03:16 PA=180 flats. Scale exptimes by 7x (so 7-sec). There are stars in these flats, which have backgrounds ~15k (g') and 10k (r').

03:23 A few PA=270 flats with exptime scaled to 10-sec.

03:30 Slew to first field (OB10)
  • John H running dowrs on this field.
    • This particular field has a strong horizontal gradient (Saturn, I assume). dowrs is not picking the pupil correctly.
  • Steve offsetting 7 arcmin to the E. (This position avoids the strong gradient, at least at g')
    • First trying dowrs again.

03:45 Now running dofpia 03:45 Starting OB
  • DOFPIA did not correct all of the astigmatism that dowrs threw in on the Blue side. +500 Z5 and +500 Z6, and +300 Z7 sent before or during 035432.
    • image jump for blue 035432.
  • Red FWHM ~ 4.4 pix = 1", but Blue IQ is not good. FWHM.
  • Backgrounds are fine in 60-sec. ~6000 in Red and 4000 in Blue.
  • Blue filter hardware failure: OB stopped with 7/18 completed (3 blue and 4 red).
2014/06/15   03:57:58.995720         OB execution error: one or more systems failed [src/lbc/actions.c:1159]
41   2014/06/15   03:57:58.876408         >>> preset/expose/save interrupted at preset time, retcode=7
40   2014/06/15   03:57:58.876307         filters system encountered a problem 13 [src/channel/channel.c:1125]
39   2014/06/15   03:57:36.235004   FILTERS    WHEEL#1    hardware failure [src/filters/filters.c:1498]
38   2014/06/15   03:57:34.904704   FILTERS    WHEEL#1    timeout error on motion [src/filters/filters.c:2239]
04:03 Playing OB10 again. Blue is running.

04:15 Repeating OB10. Blue IQ poor, Red FWHM about 4.6 pix = 1", but there is a gradient from top (best) to bottom (worst, to 5.6pix ~ 1.3").

04:23 After blue image 042225, Steve adding +500 Z5, +500 Z6 to blue.

*red image 042209 has a higher background, ~10,000 counts in 60-sec.

04:29 Steve giving -1000 Z5 and -1000 Z6 to SX. (Comparison of 042225 and 042413 showed the latter was more elongated).

04:32 2.5 cycles of OB10 were done. We'll switch to MODS.

04:32 Closing dome during switch.

05:32 Reopening, but must not point in SW (this nixes the target 1).

05:54 acqMODS J1435 (PA=-40 for HA 1.5). r 1 is slit, r 2 is field. Seeing ~ 1" on guider, 0.85" on field image.
/home/modseng/pyMODS/modsAlign/Experimental/modsAlign_v152.py -l mods1r.20140615.0001.fits mods1r.20140615.0002.fits
MODS Offset Command:
   offsetxy -3.449  5.961 rel
Manual offset of -0.2 from examination of 3.

06:08 execMODS J1435_1asec.obs (3 x 600-sec exposures - SNR is good).

06:51 acqMODS HIP71172.acq
MODS Offset Command:
   offsetxy -5.534  5.638 rel
offset 0.2" to right after checking confirmatory thru-slit image. The image was 60-sec averaging over seeing speckles. Why did the object not come up exactly at the slit center - this seems to happen more than not.

07:05 execMODS HIP71172_1asec.obs (stopped after 2 of 3).

07:05 some cirrus on all-sky, but it is wispy and well-illuminated by the moon

07:14 acqMODS bd33264_1asec.acq (not what I wanted). acqMODS bd332642.acq
  • experimental modsAlign didn't work --- not clear why or whether user error, though.
  • modsAlign
MODS Offset Command:
   offsetxy -4.108  5.421 rel

07:30 execMODS bd33...

07:43 aborted attempt to obtain spectrum through 1" slit --- the centering is not good on the acquistion image, probably a reflection of simply the jitter in 3-sec. But I have doubts how to use the 1" standard star image, so will go on to observe with LUCI.

07:46 Switching to LUCI1.
  • Detector readout has been less noisy and we'll try some of the stars in CL1806.
  • Wind speed also has dropped a little, allowing us to point to the SW.

08:05 executeLUCIScript.sh 1806_20.acq
  • seeing remains good, <~ 1".
  • with z + HeI (4 and 5), didn't see any stars, even after sky-sub. Switched to Ks (6), but object a little too bright. BrGam OK.
    • 7 is br gam field. Object is at 969.4, 1036.9
    • 8 is through slit image. Slit position is 979.5
    • offset -1.1, -9.79 calculated. This is reasonable but not exact. (0.12* diffx = -1.21)
    • 9 don't see object in slit ... offset +1.1, +9.79 to return to original position.
    • offset -1.2, 0. I still do not see the object... this is crazy - it is so bright.
    • 11, mask in turnout. object at 959.5, 1036.3
    • oh no - this was my slip! I mistook the jargon and reversed the meanings of "reference" and "new"... ugh... Actually, while the terms are confusing, I think the main problem here was that I'd just used MODS, where modsAlign asks the user to click on the slit, then target. LUCI1 "shift image" works in reverse to modsAlign...sigh...
      • guide probe failed to move, now. Tried a tiny offset along slit, but it failed. Waited for it to fail and now gave a larger offset along the slit, which this time worked.
    • 13 --- looks very well centered, finally.
    • the grating was in 2nd order --- I had used 1st order on Monday and had not thought it would have been changed since then. Should have checked, just had forgotten that there were nights of observing in between.
08:52 executeLUCIScript.sh 1806_20.spec
  • total counts in 4 dits of 30-sec each about 25000 (6000 cnts/dit). OK
  • 14/15. Unfortunately, just got the AB

08:52 light cirrus has been streaming in for about the last 15-20 minutes.

09:00 going to telluric
  • offset 1.18, -5.89

09:08 executeLUCIScript.sh Hip87881_tell.spec ---
  • seeing is puffing up to 1.5-2" now, wind has died down and cirrus is over much of the sky (not so good for the faint LBC targets, but we'll switch).

09:34 LBCs ready. Slewing to OB11 field. Needed manual Z8.
  • "S" to stop dofpia - both sides appear to have converged.

09:49 OB11 5pt dither started. There is still light cirrus, not too bad...trying the series of 5.
  • background counts are ~10,000 in both g' and r'.

09:55 Slew to OB12 field to focus/collimate.

10:03 Start OB12 - background is <12,000 in red. In Blue it is ~20,000 cnts.

10:06 Start loop.. OB13,OB14,OB12... blue filters hardware failure. Auto-recovered although we lost the blue image.

10:19 Moon-illuminated cirrus is raising the background counts, to 20,000 in r' now. The background counts are varying by a lot. After 3 cycles with 1x60sec, changed to 2x30sec, but cirrus went away. Now back to 1x60sec.
  • at 10:29 UT background on r' was about 9600 counts in 60-sec, so that is a reference. Seeing about 1.2" on guider.
  • 10:42 r' image is swamped by background, 28000 cnts.
  • 10:54 background is dominated by cirrus still, not twilight, so continuing to get a few good images. At 10:54 the r' image again has 9600 counts and we're in a clear spot.
    • cycle 1,2,3 starting at 105241 was taken in clear patch.
    • at 11:10, twilight is taking over and causing counts to increase. Stopping series.

11:14 We'll try to focus/collimate near OB11, but as it's close to moon...
  • 11:16 launching dofpia, collimated in 2 iterations --- estimated seeing is terrific: 0.77" in blue and 0.52" in red --- would that we could believe it!
  • 11:20 launching OB11_5... backgrounds all high, ~20,000. Scaled exptime by 0.5, but it is just getting too light.

11:25 Done with observing. Steve's going to test the catalog creation tool that he was working on.

11:39 Dome closed, lights off, telescope at zenith.
  • Started series of 25 LBC biases (2x). On 2nd series, blue sequence hung just before 1200 UT. No error visible in log, but at 12:28 UT, although I had stopped the OB after noticing this, I see all buttons now grey, no LST, RA, DEC, HA readings and log shows an error on blue (which may be unrelated to the image-taking stopping):
2014/06/15   12:28:48.989000   TRACKERS   CCDCTRL    safetyoff has been forced
15   2014/06/15   12:28:48.989000   TRACKERS   CCDCTRL    "FPGA not loaded" (code:48) controller error [ccdctrl.c:2526]
14   2014/06/15   12:28:47.895000   TRACKERS   CCDCTRL    safetyoff has been forced
13   2014/06/15   12:28:47.880000   TRACKERS   CCDCTRL    "FPGA not loaded" (code:48) controller error [ccdctrl.c:2520] 
  • Started MODS grlamps.cal
    • execMODS -b LS60x1.0 slitflats.cal
    • execMODS grpixflats.cal
    • execMODS bias8K.cal
    • execMODS bias1K.cal (for acq images).
  • Moved LUCI1 cal unit in
    • doing lambda cals Ar 1700 cnts in brightest line. Xe 300 cnts. This was with DIT/NDIT=5/1. Changed to settings I'd determined after some trial and error on Monday: 30x1. This gives 13000 counts in brightest line for Ar and 1600 in brightest line in Xe.
    • 34 is a test 4-sec exposure with HALO1. Subsequent exposures were taken by script, but before manually turning lamp back off. Thought script would turn lamp off to take the "darks" that are part of the lamp command. Restarting script from 0038...
    • Ran a few with Halo2 for comparison.
  • Moving LUCI cal unit out. Flushing detector for a while.

13:07
  • MODS put to sleep
  • LUCI darks started
  • LBCs turned off and disconnected.

-- %USERSIG{OlgaKuhn - 2014-06-15}%

Comments

 
Topic revision: r4 - 15 Jun 2014, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback